Object A has mass and velocity. Object B has double the mass and triple the velocity of object A. The kinetic energy of object B will be _______ times the kinetic energy of object A
1
Expert's answer
2018-05-08T10:55:08-0400
The kinematic energy by definition
KE=(mv^2)/2
Because v_B=3v_A, m_B=2m_A, we get
KE_B=((2m_A ) (3v_A )^2)/2=18 (m_A v_A^2)/2=18KE_A
Answer: 18